### v3

# AnyMaker: Zero-shot General Object Customization via Decoupled Dual-Level ID Injection

Abstract
Text-to-image based object customization, aiming to generate images with the same identity (ID) as objects of interest in accordance with text prompts and reference images, has made significant progress. However, recent customizing research is dominated by specialized tasks, such as human customization or virtual try-on, leaving a gap in general object customization. To this end, we introduce AnyMaker, an innovative zero-shot object customization framework capable of generating general objects with high ID fidelity and flexible text editability. The efficacy of AnyMaker stems from its novel general ID extraction, dual-level ID injection, and ID-aware decoupling. Specifically, the general ID extraction module extracts sufficient ID information with an ensemble of self-supervised models to tackle the diverse customization tasks for general objects. Then, to provide the diffusion UNet with the extracted ID as much while not damaging the text editability in the generation process, we design a global-local dual-level ID injection module, in which the global-level semantic ID is injected into text descriptions while the local-level ID details are injected directly into the model through newly added cross-attention modules. In addition, we propose an ID-aware decoupling module to disentangle ID-related information from non-ID elements in the extracted representations for high-fidelity generation of both identity and text descriptions. To validate our approach and boost the research of general object customization, we create the first large-scale general ID dataset, Multi-Category ID-Consistent (MC-IDC) dataset, with 315k text-image samples and 10k categories. Experiments show that AnyMaker presents remarkable performance in general object customization and outperforms specialized methods in corresponding tasks. Code and dataset will be released soon in https://github.com/LingjieKong-fdu/AnyMaker.
Figure 1: Diverse applications of AnyMaker. Given text prompts and one single reference image, our AnyMaker can achieve various customization for general objects with high ID fidelity and flexible text editability, without further fine-tuning.
1 Introduction
Recent advancements in diffusion-based text-to-image generative models [1, 2, 3] have entitled common users to create high-fidelity photo-realistic images with minimal expertise. However, generating images from specific user elements, such as a personal dog of users, remains a considerable challenge [4, 5, 6, 7] due to the difficulty of preserving the ID of specific elements. To address this limitation, object customization, which leverages reference images and textual descriptions to generate images that shares the same identity as objects of interest, has become a crucial topic.
Current object customization methods fall into two main categories: object-specific methods and object-agnostic methods. Object-specific methods, such as DreamBooth [8] and DisenBooth [9], inject customized knowledge into pretrained diffusion models via learning additional reference-specific identifiers. While these methods can be adapted to various objects, their effectiveness relies on time-consuming fine-tuning on a few reference images, which leads to the low efficiency of such methods. Object-agnostic methods, or zero-shot methods like PhotoMaker [6] and InstantID [5], on the other hand, employ large-scale training datasets to generate customized images without further fine-tuning. However, these methods predominantly concentrate on specialized tasks that are only applicable to specific domains, such as human customization or virtual try-on. In this paper, we target at zero-shot general object customization which has seldom been studied in previous works.
To this end, we propose AnyMaker, an innovative zero-shot general object customization framework in text-to-image generation tasks, which can achieve high-fidelity customization for general objects given a text prompt and one single reference image without any test-time fine-tuning. The efficacy of AnyMaker comes from its novel general ID extraction, dual-level ID injection, and ID-aware decoupling.
Specifically, in contrast to prior approaches [6, 7, 5] that merely utilize a single model to extract object representations, we employ an ensemble of self-supervised models to extract more comprehensive information from general objects, ensuring the extracted ID-related information is sufficient enough for the high-ID-fidelity customization. Then, to integrate the extracted ID information into the diffusion UNet while preserving the text editability of the model, we design a global-local dual-level ID injection strategy. At the global level, we fuse the class tokens of the extracted feature with the class word (such as dog or cat) of the text description, injecting semantic-level ID information into the diffusion UNet without damaging the other elements in the text. At the local level, we inject the patch tokens of the extracted feature into the diffusion UNet through newly added cross-attention modules without compressing the feature dimension, ensuring that the model receives a substantial amount of ID-related details, which is crucial for achieving high-ID-fidelity generation. Furthermore, we find that the ID information is usually coupled with redundant non-ID elements such as the motions and orientations of objects. Such coupling phenomena can disrupt the text editing ability and ID retention ability of the model, and further hinder the diversity of generation. To address this issue, we design a novel ID-aware decoupling module at the training stage, compelling the model to discern the intrinsic properties of the ID information and exclude non-ID information away from the generation process.
Benefiting from the innovative ID processing modules mentioned above, AnyMaker realizes zero-shot customization for general objects with high ID fidelity and flexible text editability. We demonstrate various applications of AnyMaker in Fig.1, such as realistic human customization, cartoon character customization, virtual try-on, story generation, ID mixing, and so on.
To validate our method and facilitate the development of general object customization, we build the first large-scale general ID dataset, that is Multi-Category ID-consistent (MC-IDC) dataset, which contains 315k text-image samples with over 10k categories. Experiments demonstrate that AnyMaker achieves state-of-the-art generation results on general object customization and even outperforms task-specific methods in specialized domains like human customization and virtual try-on.
In summary, the contributions of our work are as follows:
  • We propose a novel zero-shot general object customization framework in text-to-image generation tasks, which is capable of customizing general objects with high ID fidelity and flexible text editability through innovative general ID extraction, dual-level ID injection and ID-aware decoupling.
  • We construct a large-scale Multi-Category ID-Consistent (MC-IDC) dataset to significantly advance the development of general object customization.
  • We validate the effectiveness of our method through extensive experiments, which demonstrate that our AnyMaker delivers desirable performance in general object customization and surpass specialized methods tailored for specific tasks.
2 Related Work
Diffusion models. Diffusion models have demonstrated their effectiveness for text-to-image generations. The DDPM model pioneered in this field, using a diffusion and denoising process to create a mapping between Gaussian and image distributions. The Latent Diffusion Model (LDM) [10] took this a step further by applying the diffusion model to a latent space rather than pixel space, leading to the creation of text-to-image diffusion models like Stable Diffusion (SD), Midjourney, and DALLE-3 [11]. One main research problem for the diffusion model is its structure. Other than the original UNet, DiT  [12] and Pixart-$\alpha$ [3] adopt transformer as the backbone structure. Other researchers have also focused on enhancing image fidelity regarding complicated concepts. By manipulating the cross-attention maps according to text prompts,  [13] ensures that required objects are sufficiently generated in the images.  [14] proposed to process the text prompts to decompose all target attributes, which can be further used to guide the denoising procedure with the help of different objective functions. To enhance content consistency,  [15] adopted an inner localization method by calculating the cross attention. Our method is generally built on diffusion models. While different from these basic models for plain image generations, we focus on the more challenging topic of zero-shot any object customization.
Image customization. Image customization aims to generate images with the same objects of interest as the reference image [8, 16, 17, 18, 5]. Current methods can be generally grouped into two categories. One is object-specific model, which tries to finetune the pretrained diffusion models with reference images. Textual inversion [19] learns an extra word embedding for the target concept. Dreambooth [8] and DisenBooth [9] enhance such a pipeline with auxiliary objectives such as prior loss, which ensures the retention of the model’s original generative capabilities throughout the fine-tuning process. The other genre is object-agnostic models, which can directly process reference images without further fine-tuning. PhotoMaker [6], as a typical zero-shot method, is driven by training a person ID-extaction network, while  [20] uses a face model with CLIP to extract ID features. With the help of ControlNet [16], InstantID [5] can control face expression by leveraging a pre-trained face model to extract face embedding. MagicCloth [21] utilizes a ControlNet style garment extractor to encode garment details and fuse with attention for clothes replacement or preview. On the contrary to these works, we try to combine the versatility of object-specific models and the efficiency of object-agnostic models to present a method that can customize any object in a zero-shot manner, which has never been studied before.
Figure 2: Overview of our AnyMaker. AnyMaker is a zero-shot text-to-image customization method for general objects, consisting of general ID extractor, global-local dual-level ID injection, and ID-aware decoupling module.
3 Method
We first outline the overall structure of AnyMaker in Sec. 3.1. Then, we detail the general ID extraction, dual-level ID injection and ID-aware decoupling in Sec. 3.2, Sec. 3.3 and Sec. 3.4 respectively.
3.1 Overview
Give a reference image and a text prompt, our goal is to generate images that shares the same ID as the object of interest from the reference image, and modifies non-ID elements such as motions and backgrounds in accordance with the text prompt. To this end, we propose AnyMaker, an innovative zero-shot text-to-image customization framework for general objects, as summarized in Fig. 2. Firstly, we utilize the general ID extractor to acquire sufficient ID information from the reference image with a segmentation mask of the interested object. Next, we leverage the dual-level ID injection to endow the diffusion model with ID information at the global and local levels respectively, without damaging the text editing ability of the model. Additionally, we design a novel ID-aware decoupling module to distinguish the ID information from the coupled non-ID elements in the object representation, for higher ID fidelity and better text editing ability.
3.2 Genral ID Extraction
We utilize an ensemble of self-supervised models, consisting of DINOv2 [22] and MAE [23], to extract the representations of general objects in the reference image.
Previous works [6, 5] targeting at customization tasks in specific domains like human face, usually use a single model to extract object representations. However, our experiments demonstrate that one single pre-trained model, such as CLIP [24], DINOv2 [22] or MAE [23], cannot extract sufficient features for high-ID-fidelity generation of general objects, as shown in Fig. 4 and Tab. 2. Specifically, previous works [4, 5] demonstrate that features extracted by CLIP contain insufficient details, thus being less practical as an ID feature extractor, while DINOv2 is capable of providing desirable details due to its self-supervised contrastive learning strategy. However, we find that the features extracted by DINOv2 are color-insensitive because it is trained with augmentations such as ColorJitter. Consequently, solely using DINOv2 as an ID extractor would lead to mistakenly generated color. MAE, on the other hand, is aware of the color thanks to its reconstruction-related objective.
Therefore, we leverage DINOv2 and MAE, which compensate for each other, to extract sufficient and comprehensive information to handle the challenging customization task for general objects with various categories and diverse characteristics. To align the extracted features and facilitate the subsequent injection, we adopt a two-layer MLP. The process of extracting general object representations is as follows,
$f_{dino}^{C},f_{dino}^{P}=MLP(F_{dino}(I_{ref}\odot M_{ref})),$ (1)
$f_{mae}^{C},f_{mae}^{P}=MLP(F_{mae}(I_{ref}\odot M_{ref})),$ (2)
where $\odot$ represents element-wise multiplication, and $f_{dino}^{C},f_{dino}^{P},f_{mae}^{C},f_{mae}^{P}$ represent the class tokens and patch tokens extracted by DINOv2 and MAE respectively. $F_{dino}$ and $F_{mae}$ represent the DINOv2 and MAE backbone. $I_{ref}$ represents the reference image, and $M_{ref}$ is the segmentation mask of the interested object in the reference image. Consequently, tokens derived from our general ID extractor not only offer intricate details from DINOv2, but also have sufficient colour and structural information from MAE.
3.3 Dual-level ID Injection
We propose a dual-level ID injection mechanism to inject the global semantic ID information and local fine-grained ID information into the diffusion UNet respectively, ensuring both flexible text editability and high ID fidelity in the customization process for general objects.
Global ID injection. In text-to-image generation, text prompts are leveraged to guide diffusion models to generate diverse images. Nevertheless, diffusion models are incapable of accurately generating images with a specific object ID solely relying on the text description. To inject the ID information of the interested objects into the text, while simultaneously maintaining the editing capability of text prompts for other elements such as the background and motions, we design the global ID injection mechanism. Specifically, we merge the semantic-level class tokens $(f_{dino}^{C},f_{mae}^{C})$ with the class word in the text embedding, such as "dog" or "cat",
$f_{fuse}^{C}=MLP(Concat(f_{text}^{C},f_{dino}^{C},f_{mae}^{C})),$ (3)
where $f_{fuse}^{C}$ is the fused class token, and $f_{text}^{C}$ represents the class word embedding of the interested object. Then we insert the fused class token $f_{fuse}^{C}$ to the original position of the class word in the text embeddings, obtaining the global-level condition $c_{g}$ with ID-related information of the object of interest, as shown in Fig. 2. Following this, the global-level condition $c_{g}$ engages with the cross-attention module in the UNet, like to standard text-to-image models [10], as follows,
$Z_{g}=Attention(Q_{g},K_{g},V_{g}),$ (4)
where $Q_{g}=Z_{g}W_{q}$, $K_{g}=c_{g}W_{k}$, $V_{g}=c_{g}W_{v}$ are the query, key, and value of cross-attention module, in which $W_{q}$, $W_{k}$, $W_{v}$ are the weight matrices of the trainable linear projection layers, and $Z_{g}$ is the query feature containing the latent input information.
Local ID injection. Semantic-level information in the global-level condition $c_{g}$ contains relatively fewer details of the objects, and thus the model cannot handle complex customization tasks for general objects effectively merely relying on global ID injection. To supply the model with sufficient ID-related details, we propose the local ID injection mechanism. Specifically, we first fuse the extracted patch tokens $f_{dino}^{P}$ and $f_{mae}^{P}$ with a two-layer MLP to densify the contained information to a unified representation $c_{l}$, as follows,
$c_{l}=MLP(f_{dino}^{P})+MLP(f_{mae}^{P}),$ (5)
where $c_{l}$ serves as the local-level condition for the diffusion UNet. Then, we add one cross-attention module in each upblock of the diffusion UNet in order to inject the ID-related details into the model without compressing $c_{l}$ dimensions, feeding the model with as much ID information as possible.
$Z_{l}=Attention(Q_{l},K_{l},V_{l}),$ (6)
where $Q_{l}=Z_{l}W_{q}$, $K_{l}=c_{l}W_{k}$, $V_{l}=c_{l}W_{v}$ are the query, key, and value of cross-attention module, in which $W_{l}$, $W_{l}$, $W_{l}$ are the weight matrices of the trainable linear projection layers, and $Z_{l}$ is the query feature containing the latent input information.
Through the dual-level ID injection, we provide the model with sufficient ID-related information, both at the semantic level and the detail level, on the premise of not impairing the text editing ability of the model.
3.4 Training with ID-Aware Decoupling
The ID-related information injected in the diffusion model is usually entangled with redundant non-ID information such as motion, direction and size. Without proper guidance, the model would learn both kinds of information together, which may disrupt the text editing ability of the model and hinder the diversity of generation. For instance, in a reference image depicting "a standing person", the posture "standing" may also be extracted and injected into the diffusion model, which could impede a text prompt wishing to generate "a sitting person". Consequently, the generated images would closely resemble the provided reference images, thereby compromising the text editing ability and generation diversity.
To eliminate the negative impact brought by the coupling phenomenon, we take inspiration from previous works [9, 23] to propose an ID-aware decoupling module, which consists of a decoupling branch and a group of losses based on the normal generation branch at the training stage, guiding the model to discern the intrinsic properties of ID information.
Decoupling branch and normal branch. In the decoupling branch, we first extract the target image features $f_{tar}$ through CLIP [24] to make an image embedding prior [11]. Next, we mask out the ID information contained in the image embedding by a trainable feature mask $m_{id}$, and thus the masked image feature $f_{msk}=f_{tar}\odot m_{id}$ contains only non-ID information of the target image, where $\odot$ represents element-wise multiplication. Then, we add the masked image feature $f_{msk}$ to the fused class token $f_{fuse}^{C}$ and then feed it to the diffusion UNet for generation following the injection way mentioned in the Sec. 3.3. Yet in the normal branch, we inject $f_{fuse}^{C}$ to the model without $f_{msk}$. The generation process of two branches can be illustrated as follows,
$f_{msk}\oplus f_{fuse}^{C}\Rightarrow I_{tar},\hskip 14.22636ptf_{fuse}^{C}\Rightarrow I_{tar},\hskip 14.22636ptf_{msk}\perp f_{fuse}^{C},$ (7)
where $\oplus$ represents element-wise addition and $I_{tar}$ is the target image, and the symbol $\perp$ signifies that the features it connects are independent of each other and have no obvious correlation. Through training in the two branches, we push the non-ID information into the masked image feature $f_{msk}$ and only retain the ID-related information in the fused class token $f_{fuse}^{C}$.
Training strategy. We use three losses in our training stage to ensure that the ID-aware decoupling module takes effect. Specifically, We utilize the following denoising losses to train each branch,
$\mathcal{L}_{decouple}=\|\epsilon-\epsilon_{\theta}\big{(}x_{t},c_{g},f_{msk},c_{l},t\big{)}\|^{2},$ (8)
$\mathcal{L}_{normal}=\|\epsilon-\epsilon_{\theta}\big{(}x_{t},c_{g},c_{l},t\big{)}\|^{2},$ (9)
where $t$ is the randomly sampled time step and $x_{t}$ represents the noisy latent of the target image, $\epsilon$ is the ground truth noise and $\epsilon_{\theta}$ is the predicted noise. $c_{g}$ and $c_{l}$ represent the global-level and local-level conditions acquired in Sec. 3.3. Additionally, we design a contrastive loss, ensuring that the masked image feature $f_{msk}$ captures non-ID information and the fused class token $f_{fuse}^{C}$ remains free of it, as below,
$\mathcal{L}_{contrast}=Sim(f_{fuse}^{C},f_{msk}),$ (10)
where $Sim$ is instantiated as cosine similarity. In summary, the training loss is the sum of the three loss functions mentioned above,
$\mathcal{L}=\alpha_{1}\mathcal{L}_{normal}+\alpha_{2}\mathcal{L}_{decouple}+\alpha_{3}\mathcal{L}_{contrast},$ (11)
where $\alpha_{1},\alpha_{2},\alpha_{3}$ represent hyperparameters to balance the three losses.
4 MC-IDC Dataset
We create the first large-scale general ID dataset, named as Multi-Category ID-Consistent (MC-IDC) dataset, designed for the research of general object customization. We introduce the characteristics and construction process of the dataset in Sec. 4.1 and Sec. 4.2 respectively. We introduce more details of our dataset in the Appendix.
4.1 Characteristics
Large-scale dataset. Our dataset consists of approximately 315,000 samples in total with more than 10,000 categories, covering various types such as human faces, animals, clothes, human-made tools, etc. Adequate training samples and diverse categories endow the model with general ID understanding. We introduce the main categories of the dataset and their quantities in the Appendix.
High quality. The average resolution of the images in our dataset is 1039$\times$950, which is helpful for the high-quality generation of our model in both training and inference stage.
Reference-target image pair. Each sample consists of a reference image, a segmentation mask of the object of interest in the reference image, a target image, and a text caption of the target image. During training, the reference image with its segmentation mask provides the model with ID information, the text caption of the target image offers semantic-level guidance for generation, and the target image serves as the ground truth. In each sample, the reference image and the target image depict the same object in different states, which means that each reference-target image pair shares the same ID of the object of interest and has different non-ID elements such as motions and directions. Training with reference-target image pairs endows the model with the ID processing ability while not damaging the diversity of generation.
4.2 Construction Pipeline
The construction pipeline of MC-IDC consists of four steps: data collection, instance detection and segmentation, image pair generation, and text prompt generation.
Data collection. MC-IDC dataset contains diverse sources including web-crawled images, movies and publicly available datasets. To facilitate the subsequent generation of image pairs, most of the data we collect are video datas or multi-view datas. To ensure the high quality of our dataset, we delete the images with a resolution less than 300$\times$300.
Instance detection and segmentation. For some publicly available datasets or web-crawled images which do not provide segmentation annotations, we use the advanced instance segmentation model [25] for instance detection and segmentation. For movie dataset, we extract the frames with 10fps and do object tracking with  [26], aiming to establish the ID connection across frames.
Image pair generation. For most data samples which are from a clip of video or a group of multi-view data, we randomly select two frames containing the same object and perform a random crop around the object to form the reference-target image pair. For the other data samples from the single image dataset, we perform random data augmentations on the single image two times, and take them as the reference image and the target image respectively.
Text prompt generation. In each image pair, the target image is densely captioned with the state-of-the-art large vision language model [27], serving as the text prompt to guide the generation.
For more details about the MC-IDC, please refer to the Appendix.
Table 1: Quantitative comparison among zero-shot object customization methods in general domain and two specific popular domains namely human customization and virtual try-on.
[…]
General objects IP-Adapter 70.32 77.18 28.03 44.94 - 84.43±0.66
Ours 47.09 82.16 29.27 65.13 - 74.38±3.99
Human customization IP-Adapter 102.69 72.17 29.32 41.44 65.51 -
PhotoMaker 106.35 71.80 32.13 44.62 64.10 -
InstanceID 113.18 75.87 32.89 49.26 63.26 -
Ours 86.40 79.60 30.88 57.44 78.54 -
Virtual try-on MagicClothing 126.09 76.53 21.40 29.10 - 89.36±0.40
IP-Adapter 104.47 81.99 25.03 59.39 - 71.28±5.06
Ours 50.65 83.82 22.42 66.24 - 71.27±3.63
5 Experiment
5.1 Setup
Implementation details. To be compatible with open-community as Civitai.com, we employ SD1.5 [28], the most used foundational model in the public, as the backbone model in our experiments. We use the multi-scale training mode proposed in [1] considering the resolutions of images from different sources are diverse. We employ the same CLIP image encoder as in [28] in the ID-aware decoupling module. We use MAE [23] and DINOv2 [22] for extracting ID features. During training, we set the learning rate as 1e-5 with a batch size of 32. We conduct the training process about 6 epochs with 32 V100 GPUs, consuming about 30 hours. We adopt dropout with a ratio of 0.1 for the global-level condition $c_{g}$ and the local-level condition $c_{l}$ respectively. During inference, we utilize 50 steps denoising process for generation and set the classifier-free guidance scale as 7.
Competitors. To demonstrate the generality of our AnyMaker, we not only make a comparison with previous works on general object customization, but also compare it with task-specialized methods in two popular specific customization domains, that is, human customization and virtual try-on. Specifically, we compare AnyMaker with IP-Adapter [7] on zero-shot general object customization. On human customization, we further choose human specialized method PhotoMaker [6] and InstantID [5] for comparison. On virtual try-on, we conduct a comparison with the recent state-of-the-art MagicClothing  [21]. It is worth noting that all the methods mentioned above are zero-shot methods and are capable of customizing images based solely on one single reference image. Therefore, in this paper, we do not compare object-specific methods such as DreamBooth [8] and DisenBooth [9] which typically require 3 to 5 reference images and additional fine-tuning.
Evaluation dataset. Our evaluation dataset contains 1000 text-image samples, including general objects, human data, and virtual try-on data with a ratio of 4:3:3. All samples in the evaluation dataset are randomly selected from the MC-IDC and do not appear during the training stage.
Evaluation metrics. Following PhotoMaker [6], we utilize DINO-i [22] and CLIP-i [29] to measure the ID fidelity and use CLIP-t to measure the prompt fidelity. We leverage FID [30] to assess the generation quality. For human customization, we additionally calculate the face similarity (FaceSim) with FaceNet [31] as commonly done in [6, 5]. Furthermore, we design a novel metric DiverSim-i, which calculates the average DINO similarity among the images generated according to diverse text prompts containing different scenarios. The lower the value of DiverSim-i, the stronger the ability of the model to generate diverse images in line with text prompts.
We introduce more details about the experimental setup in the Appendix.
Figure 3: Qualitative comparison on zero-shot general object customization and two specific popular tasks namely human customization and virtual try-on. The AnyMaker exhibits great ID-preserving ability with better text controls and more diverse generations on both general objects and specialized domains.
5.2 Quantative Analysis
Our AnyMaker outperforms previous works on general object customization for all metrics, and achieve comparable or better performance in contrast with specific methods tailored for specialized tasks such as human customization and virtual try-on, as shown in Tab. 1. Specifically, we achieve higher ID fidelity for CLIP-i and DINO-i, better prompt fidelity for CLIP-t, and higher generation quality evaluated by FID. Furthermore, our method exhibits better diversity on various scenarios as presented by DiverSim-i.
Table 2: Ablation study on the ID extraction methods.
Methods FID $\downarrow$ CLIP-i $\uparrow$ DINO-i $\uparrow$
CLIP 49.11 79.58 59.45
DINO 48.89 80.82 63.71
MAE 49.00 79.09 59.79
Ours 47.50 81.86 65.12
Table 3: Ablation study on the ID injection methods.
Methods FID $\downarrow$ CLIP-i $\uparrow$ DINO-i $\uparrow$
Global 49.78 78.76 60.67
Local 48.66 81.24 62.89
Ours 47.50 81.86 65.12
Table 4: Ablation study on the ID-aware decoupling module.
Methods FID $\downarrow$ CLIP-i $\uparrow$ DINO-i $\uparrow$
w.o. decoupling 47.50 81.86 65.12
with decoupling 47.09 82.16 65.13
5.3 Qualitative Analysis
The AnyMaker exhibits outstanding capabilities of high-quality customization for general objects, and even beat task-specialized methods in the specific domains, such as human customization and virtual try-on, as shown in Fig. 3. Our method is capable of generating high-ID-fidelity images given one single reference image, and simultaneously editing non-ID elements like posture or background in accordance with text prompts, which endows the model with the ability to generate diverse images, such as various scenarios and different postures. Furthermore, benefiting from the ID-aware decoupling module, our model is capable of maintaining the ID fidelity of objects while diversifying the non-ID elements such as motions and directions compared to the reference images, even without corresponding guidance of text prompts.
Figure 4: Ablation study on different ID extraction methods in the quantitative perspective. Our ID extractor has both the rich details from DINOv2 and the color information from MAE.
Figure 5: Ablation study on whether to add the ID-aware decoupling module in the quantitative perspective. The model with the ID-aware decoupling module demonstrates stronger text control ability, as shown by the blue text.
5.4 Ablation Study
Effectiveness of general ID Extraction. We compare our proposed general ID extractor with DINOv2 [22], MAE [23] and CLIP [29]. As in Tab. 2, our method performs the best, whether in terms of FID, which measures the quality of generation, or in terms of CLIP-i and DINO-i, which measure ID fidelity. As in Fig. 4, merely using DINOv2 as the extractor fails to maintain color consistency, while solely using MAE as the extractor lacks sufficient details. In contrast, our method combines the advantages of both DINOv2 and MAE, which enables our model to customize images with rich details and the consistent color of the object in reference images.
Benefits of dual-level ID injection. We explore the effectiveness of our dual-level ID injection module compared to only global-level injection and only local-level injection separately. As shown in Tab. 3, by combining global and local injection, AnyMaker achieves the best generation results both in quality as measured by FID and ID-fidelity as measured by CLIP-i and DINO-i.
Benefits of ID-aware decoupling. We conduct a comparative experiment on whether to add the ID decoupling module to the model during training, aiming to verify its effectiveness. As in Tab.4, the model trained with the ID-aware decoupling module achieves higher ID-fidelity scores in terms of CLIP-i and DINO-i, which means that the decoupling module can help the model to discern ID information embedded in object representations, thereby generating results with better ID consistency. Further, we visualize the results by models trained with and without decoupling in Fig.5. The model trained with decoupling exhibit enhanced capabilities in distinguishing ID information from non-ID elements such as motions (open or shut the mouth) and directions (front-facing or side-facing) of objects of interest. The enhanced discrimination ability allows the model to mitigate the influence of non-ID information during generation, thereby better preserving the text editing capabilities.
6 Conclusion
We introduce AnyMaker, a zero-shot text-to-image framework for general object customization, which has seldom been studied in previous works. Our approach integrates novel general ID extraction, dual-level ID injection, and ID-aware decoupling module, endowing the model with high ID fidelity without damaging the text editing ability. Experiments demonstrate that our method excels in general object customization solely with one single reference image, and even outperforms task-specialized methods in some popular specific domains. Furthermore, we construct the first large-scale general ID dataset, MC-IDC, to set a new benchmark and promote the future research for object customization.
Appendix A Appendix
A.1 Additional Details about MC-IDC
Data sources. The data sources of MC-IDC can be divided into three categories: public datasets, web-crawled images, and movies. We detail the statistical information about various data sources in Tab. A1.
Main categories. We record several main categories that appear most frequently in MC-IDC, as shown in Tab. A2.
Figure A1: Illustration of our general ID dataset, MC-IDC. In each sample, the reference image with the object mask provides ID information, the text prompt offers semantic-level guidance for generation, and the target image serves as the ground truth.
Dataset illustration. We present the overall illustration of the MC-IDC as shown in Fig. A1.
Table A1: Details about data sources of MC-IDC.
Source Dataset Type Image pair numbers
Public datasets HumanFace [32] Video 55,830
VOS [33] Video 55,823
VIPSEG [34] Video 27,983
MVImgNet [35] Multi-view image 53,909
VITON [36] Multi-view image 20,000
LVIS [37] Single image 8,003
Web-crawled images - Single image 55,829
Movies - Video 38,405
Table A2: Sample numbers of main categories in MC-IDC.
Categories Images
man 46,720
woman 26,670
clothes 20,040
girl 3,498
panda 3,007
train 2,286
car 1,974
boy 1,855
dog 1,820
A.2 Additional Details about Experiment Setup
Categories in evaluation dataset. The evaluation dataset can be divided into general objects, human data, and virtual try-on data. The human data and the virtual try-on data each contain 300 different samples. General objects in the evaluation dataset consist of 50 categories, each of which contains 8 diverse samples. We summarize the 50 categories in Tab. A3.
Table A3: Categories of general objects in the evaluation dataset.
Winter melon Cabbage Vessel Pillow Screw driver
Pants Computer mouse Lipstick Rice cooker Toy figure
Clothing Pineapple Can Plush toy Grape
Toilet paper Paper box Skirt Pawpaw Ginger
Bowl Train Bottle Cantaloupe Sanitary napkin
Soccer Bag Umbrella Hammer Book
Flower Shoe Towel Ashcan Telephone
Faucet Flowerpot Motorcycle Mug Kiwi
Pot Grapefruit Jug Car Basket
Balloons Tomato Flashlight Bagged snacks Toy duck
Text prompts for calculating DiverSim-i. We use diverse text prompts describing different scenarios to guide the generation, and calculate DiverSim-i among the generated images. We record the text prompts in Tab. A4.
Table A4: Text prompts for calculating DiverSim-i.
Scenarios Text prompts
Snow Original text prompt + "The scene of the picture is in the snow."
Original text prompt + "The background of the picture is in the snow."
Grass Original text prompt + "The scene of the picture is on the grass."
Original text prompt + "The background of the picture is on the grass."
Beach Original text prompt + "The scene of the picture is on the beach."
Original text prompt + "The background of the picture is on the beach."
Jungle Original text prompt + "The scene of the picture is in the jungle."
Original text prompt + "The background of the picture is in the jungle."
Eiffel Tower Original text prompt + "The scene of the picture is beside the Eiffel Tower."
Original text prompt + "The background of the picture is beside the Eiffel Tower."
A.3 More Visual Results
Our AnyMaker exhibits outstanding performance on various applications, such as general object customization in Fig. A2, virtual try-on in Fig. A3, Text-Image ID Mixing in Fig. A4, ID-consistent generation in Fig. A5, and story generation in Fig. A6
Figure A2: Additional visual results of general object customization. Our AnyMaker is capable of generating high-ID-fidelity images given one single reference image, and simultaneously editing non-ID elements like posture or background in accordance with text prompts.
Figure A3: Additional visual results of virtual try-on. Given a piece of clothing, the AnyMaker can generate images of the clothing worn on a person.
Figure A4: Additional visual results of text-image ID mixing. If the category of the interested object in the text prompt and that in the reference image is not the same, our AnyMaker can merge the two and form a new ID.
Figure A5: Additional visual results of ID-consistent generation. The AnyMaker can generate multiple ID-consistent images with diverse non-ID elements such as motions and orientations.
Figure A6: Additional visual results of story generation. Our AnyMaker can generate diverse images under the guidance of text prompts, while maintaining the same identity as the object of interest in the reference image, thereby enabling the creation of a cohesive narrative.

### v4

@@
-AnyMaker: Zero-shot General Object Customization via Decoupled Dual-Level ID Injection
+CustAny: Customizing Anything from A Single Example
 Abstract
-Text-to-image based object customization, aiming to generate images with the same identity (ID) as objects of interest in accordance with text prompts and reference images, has made significant progress.
-However, recent customizing research is dominated by specialized tasks, such as human customization or virtual try-on, leaving a gap in general object customization.
-To this end, we introduce AnyMaker, an innovative zero-shot object customization framework capable of generating general objects with high ID fidelity and flexible text editability.
-The efficacy of AnyMaker stems from its novel general ID extraction, dual-level ID injection, and ID-aware decoupling.
-Specifically, the general ID extraction module extracts sufficient ID information with an ensemble of self-supervised models to tackle the diverse customization tasks for general objects.
-Then, to provide the diffusion UNet with the extracted ID as much while not damaging the text editability in the generation process, we design a global-local dual-level ID injection module, in which the global-level semantic ID is injected into text descriptions while the local-level ID details are injected directly into the model through newly added cross-attention modules.
-In addition, we propose an ID-aware decoupling module to disentangle ID-related information from non-ID elements in the extracted representations for high-fidelity generation of both identity and text descriptions.
-To validate our approach and boost the research of general object customization, we create the first large-scale general ID dataset, Multi-Category ID-Consistent (MC-IDC) dataset, with 315k text-image samples and 10k categories.
-Experiments show that AnyMaker presents remarkable performance in general object customization and outperforms specialized methods in corresponding tasks.
-Code and dataset will be released soon in https://github.com/LingjieKong-fdu/AnyMaker.
-Figure 1: Diverse applications of AnyMaker.
-Given text prompts and one single reference image, our AnyMaker can achieve various customization for general objects with high ID fidelity and flexible text editability, without further fine-tuning.
+Recent advances in diffusion-based text-to-image models have simplified creating high-fidelity images, but preserving the identity (ID) of specific elements, like a personal dog, is still challenging.
+Object customization, using reference images and textual descriptions, is key to addressing this issue.
+Current object customization methods are either object-specific, requiring extensive fine-tuning, or object-agnostic, offering zero-shot customization but limited to specialized domains.
+The primary issue of promoting zero-shot object customization from specific domains to the general domain is to establish a large-scale general ID dataset for model pre-training, which is time-consuming and labor-intensive.
+In this paper, we propose a novel pipeline to construct a large dataset of general objects and build the Multi-Category ID-Consistent (MC-IDC) dataset, featuring 315k text-image samples across 10k categories.
+With the help of MC-IDC, we introduce Customizing Anything (CustAny), a zero-shot framework that maintains ID fidelity and supports flexible text editing for general objects.
+CustAny features three key components: a general ID extraction module, a dual-level ID injection module, and an ID-aware decoupling module, allowing it to customize any object from a single reference image and text prompt.
+Experiments demonstrate that CustAny outperforms existing methods in both general object customization and specialized domains like human customization and virtual try-on.
+Our contributions include a large-scale dataset, the CustAny framework and novel ID processing to advance this field.
+Code and dataset will be released soon in https://github.com/LingjieKong-fdu/CustAny.
 1 Introduction
-Recent advancements in diffusion-based text-to-image generative models [1, 2, 3] have entitled common users to create high-fidelity photo-realistic images with minimal expertise.
-However, generating images from specific user elements, such as a personal dog of users, remains a considerable challenge [4, 5, 6, 7] due to the difficulty of preserving the ID of specific elements.
-To address this limitation, object customization, which leverages reference images and textual descriptions to generate images that shares the same identity as objects of interest, has become a crucial topic.
-Current object customization methods fall into two main categories: object-specific methods and object-agnostic methods.
-Object-specific methods, such as DreamBooth [8] and DisenBooth [9], inject customized knowledge into pretrained diffusion models via learning additional reference-specific identifiers.
-While these methods can be adapted to various objects, their effectiveness relies on time-consuming fine-tuning on a few reference images, which leads to the low efficiency of such methods.
-Object-agnostic methods, or zero-shot methods like PhotoMaker [6] and InstantID [5], on the other hand, employ large-scale training datasets to generate customized images without further fine-tuning.
-However, these methods predominantly concentrate on specialized tasks that are only applicable to specific domains, such as human customization or virtual try-on.
-In this paper, we target at zero-shot general object customization which has seldom been studied in previous works.
-To this end, we propose AnyMaker, an innovative zero-shot general object customization framework in text-to-image generation tasks, which can achieve high-fidelity customization for general objects given a text prompt and one single reference image without any test-time fine-tuning.
-The efficacy of AnyMaker comes from its novel general ID extraction, dual-level ID injection, and ID-aware decoupling.
-Specifically, in contrast to prior approaches [6, 7, 5] that merely utilize a single model to extract object representations, we employ an ensemble of self-supervised models to extract more comprehensive information from general objects, ensuring the extracted ID-related information is sufficient enough for the high-ID-fidelity customization.
-Then, to integrate the extracted ID information into the diffusion UNet while preserving the text editability of the model, we design a global-local dual-level ID injection strategy.
-At the global level, we fuse the class tokens of the extracted feature with the class word (such as dog or cat) of the text description, injecting semantic-level ID information into the diffusion UNet without damaging the other elements in the text.
-At the local level, we inject the patch tokens of the extracted feature into the diffusion UNet through newly added cross-attention modules without compressing the feature dimension, ensuring that the model receives a substantial amount of ID-related details, which is crucial for achieving high-ID-fidelity generation.
-Furthermore, we find that the ID information is usually coupled with redundant non-ID elements such as the motions and orientations of objects.
-Such coupling phenomena can disrupt the text editing ability and ID retention ability of the model, and further hinder the diversity of generation.
-To address this issue, we design a novel ID-aware decoupling module at the training stage, compelling the model to discern the intrinsic properties of the ID information and exclude non-ID information away from the generation process.
-Benefiting from the innovative ID processing modules mentioned above, AnyMaker realizes zero-shot customization for general objects with high ID fidelity and flexible text editability.
-We demonstrate various applications of AnyMaker in Fig.1, such as realistic human customization, cartoon character customization, virtual try-on, story generation, ID mixing, and so on.
-To validate our method and facilitate the development of general object customization, we build the first large-scale general ID dataset, that is Multi-Category ID-consistent (MC-IDC) dataset, which contains 315k text-image samples with over 10k categories.
-Experiments demonstrate that AnyMaker achieves state-of-the-art generation results on general object customization and even outperforms task-specific methods in specialized domains like human customization and virtual try-on.
-In summary, the contributions of our work are as follows:
-  • We propose a novel zero-shot general object customization framework in text-to-image generation tasks, which is capable of customizing general objects with high ID fidelity and flexible text editability through innovative general ID extraction, dual-level ID injection and ID-aware decoupling.
-  • We construct a large-scale Multi-Category ID-Consistent (MC-IDC) dataset to significantly advance the development of general object customization.
-  • We validate the effectiveness of our method through extensive experiments, which demonstrate that our AnyMaker delivers desirable performance in general object customization and surpass specialized methods tailored for specific tasks.
+Recent advancements in diffusion-based text-to-image generative models (e.g., SDXL, Improving GAN, PixArt Alpha) [21, 3, 6] have empowered users to generate highly realistic images with minimal technical knowledge.
+However, generating images of specific objects [8, 30, 14, 33], such as personal pets, remains challenging due to the difficulty of preserving their unique identity.
+Object customization, a technique that utilizes reference images and textual descriptions to generate images of specific objects, has emerged as a critical area of research.
+Existing methods [27, 35, 20, 18, 30] for object customization, such as DreamBooth, PortraitBooth and InstantID, have limitations in general-purpose object customization: Object-specific models [27, 5], while powerful, are often inefficient due to the need for extensive fine-tuning for each new object.
+Conversely, object-agnostic models [14, 30] have limited capabilities and often struggle with diverse object types.
+Unlike existing customizing applications tailored to one specific task like face generation or virtual try-ons, this paper focuses on general object customization from a single example, as shown in Fig.
+LABEL:fig:intro.
+Specifically, it addresses zero-shot object customization in various scenarios where no fine-tuning on specific objects is available.
+Achieving this first requires a large-scale dataset containing various categories of objects for model pre-training.
+At present, however, there are few publicly released datasets in the field of zero-shot customization for general objects.
+To address this issue, we propose a novel construction pipeline for general ID dataset and introduce the Multi-Category ID-Consistent (MC-IDC) dataset, the first large-scale dataset for zero-shot object customization across diverse scenarios.
+It includes over 315K high-quality images across 10K categories, such as human faces, animals, clothing, and tools, supporting a wide range of object types and domains.
+The dataset features reference-target image pairs with consistent object IDs, including segmentation masks and text captions, making it ideal for ID-consistent generation tasks and advancing research in general object customization.
+Based on our MC-IDC, we design a zero-shot general object customization framework which is able to Customize Anything (CustAny) from a single example.
+The innovation of our framework is mainly reflected in managing and utilizing identity information beyond the narrow focus of previous work.
+Specifically, the CustAny framework follows three essential steps: ID extraction, ID injection, and ID disentanglement.
+These steps address three critical questions: (1) How to capture enough identity details to handle a variety of complex customization tasks?
+(2) How to add this identity information to the model while still enabling text-based edits?
+(3) How to reduce the impact of irrelevant features from the reference object on the final results?
+Formally, CustAny features an ID extraction module, dual-level ID injection, and an ID-aware decoupling module, offering innovative ID handling methods.
+Unlike prior methods [14, 33, 30] that use a single model for object representations, we leverage multiple pre-trained self-supervised models to capture more detailed ID-related features for high-ID-fidelity customization.
+To integrate this information into the diffusion model, we use a dual-level ID injection strategy: globally, by merging semantic-level ID features with text descriptions, and locally, by injecting patch tokens through cross-attention modules, preserving both ID fidelity and text editability.
+Furthermore, we identify that ID information is often mixed with redundant non-ID elements, like object motion and orientation, which can disrupt both text editing and ID retention.
+To solve this, we introduce an ID-aware decoupling module during training to help the model separate ID-related properties from non-ID elements, enhancing both ID fidelity and generation diversity.
+With these advanced ID processing modules, CustAny enables zero-shot customization for various objects while supporting flexible text-based editing.
+As shown in Fig.
+LABEL:fig:intro, CustAny applies to object customization in various scenarios such as human customization, cartoon character personalization, virtual try-on.
+We conduct extensive experiments showing that CustAny achieves state-of-the-art performance in general object customization and outperforms task-specific methods in areas like human customization and virtual try-on.
+In summary, we have these contributions: (1) Dataset: We propose a novel construction pipeline for general ID dataset, and present a large-scale dataset, MC-IDC, with 315K images across 10K categories for zero-shot object customization.
+(2) CustAny Framework: We design a zero-shot framework for general object customization in text-to-image generation, focusing on ID fidelity and text-based editing.
+(3) ID Processing: CustAny uses multiple self-supervised models and a dual-level ID injection strategy to integrate ID information while preserving text-editability.
+We present a module to separate ID from non-ID elements, improving ID fidelity and generation diversity.
+(4) Good Results: CustAny outperforms task-specific methods in general and specialized object customization tasks, including human customization and virtual try-on.
+Figure 1: Illustration of our general ID dataset, MC-IDC.
+In each sample, the reference image with the object mask provides ID information, the text prompt offers semantic-level guidance for generation, and the target image serves as the ground truth.
 2 Related Work
@@
 [9] proposed to process the text prompts to decompose all target attributes, which can be further used to guide the denoising procedure with the help of different objective functions.
-To enhance content consistency,  [15] adopted an inner localization method by calculating the cross attention.
+To enhance content consistency, Tewel et al.
+[29] adopted an inner localization method by calculating the cross attention.
 Our method is generally built on diffusion models.
@@
 Image customization.
-Image customization aims to generate images with the same objects of interest as the reference image [8, 16, 17, 18, 5].
-Current methods can be generally grouped into two categories.
-One is object-specific model, which tries to finetune the pretrained diffusion models with reference images.
-Textual inversion [19] learns an extra word embedding for the target concept.
-Dreambooth [8] and DisenBooth [9] enhance such a pipeline with auxiliary objectives such as prior loss, which ensures the retention of the model’s original generative capabilities throughout the fine-tuning process.
-The other genre is object-agnostic models, which can directly process reference images without further fine-tuning.
-PhotoMaker [6], as a typical zero-shot method, is driven by training a person ID-extaction network, while  [20] uses a face model with CLIP to extract ID features.
-With the help of ControlNet [16], InstantID [5] can control face expression by leveraging a pre-trained face model to extract face embedding.
-MagicCloth [21] utilizes a ControlNet style garment extractor to encode garment details and fuse with attention for clothes replacement or preview.
-On the contrary to these works, we try to combine the versatility of object-specific models and the efficiency of object-agnostic models to present a method that can customize any object in a zero-shot manner, which has never been studied before.
-Figure 2: Overview of our AnyMaker.
-AnyMaker is a zero-shot text-to-image customization method for general objects, consisting of general ID extractor, global-local dual-level ID injection, and ID-aware decoupling module.
+Object customization techniques generally fall into two categories: object-specific and object-agnostic (zero-shot) methods.
+Object-specific approaches, like DreamBooth [27] and DisenBooth [5], fine-tune pretrained diffusion models by incorporating new, object-specific identifiers from a small set of reference images.
+While effective, these methods require extensive fine-tuning for each new object, which limits their scalability and efficiency.
+In contrast, object-agnostic or zero-shot approaches such as PhotoMaker [14] and InstantID [30], leverage large-scale training datasets to customize images without the need for additional fine-tuning.
+Although more efficient, these methods are often domain-specific, focusing on tasks like human customization, facial manipulation, or virtual try-on applications.
+For example, PhotoMaker trains a person ID extraction network, while InstantID applies ControlNet [35] to control facial expressions based on pretrained embeddings, and MagicCloth [7] uses ControlNet to customize garment details for virtual try-on.
+These existing methods, however, have limitations when it comes to general-purpose object customization.
+Object-specific models are powerful but inefficient, requiring time-intensive training for each new item, while object-agnostic models tend to be narrow in focus and are often limited to specific types of objects or applications.
+In this work, we aim to bridge these gaps by combining the versatility of object-specific methods with the efficiency of zero-shot methods.
+As illustrated in Fig.
+LABEL:fig:intro, our approach targets the ability to customize any object based on a single example image and text description, achieving generalization across diverse categories with minimal fine-tuning.
+This design marks a significant step toward general object customization across various domains, an area that remains largely unexplored.
+Figure 2: Overview of our CustAny.
+CustAny is a zero-shot text-to-image customization method for general objects, consisting of general ID extractor, global-local dual-level ID injection, and ID-aware decoupling module.
+3 MC-IDC Dataset
+We curate the Multi-Category ID-Consistent (MC-IDC) dataset, the first large-scale dataset for general object customization research.
+We present the overall illustration of the MC-IDC as shown in Fig.
+1.
+The dataset’s merits and construction process are described below, with additional details in Supplementary.
+Construction pipeline.
+The MC-IDC is constructed by the following four steps: (1) Data collection.
+MC-IDC dataset contains diverse sources including web-crawled images, movies and publicly available datasets.
+To facilitate the subsequent generation of image pairs, most of the data we collect are video datas or multi-view datas.
+To ensure the high quality of our dataset, we delete the images with a resolution less than 300$\times$300.
+(2) Instance detection and segmentation.
+For some publicly available datasets or web-crawled images which do not provide segmentation annotations, we use advanced instance segmentation model [31] for instance detection and segmentation.
+For movie dataset, we extract frames with 10fps and do object tracking with [19], aiming to establish the ID connection across frames.
+(3) Image pair generation.
+For most data samples which are from a clip of video or a group of multi-view data, we randomly select two frames containing the same object and perform a random crop around the object to form the reference-target image pair.
+For the other data samples from the single image dataset, we perform random data augmentations on the single image two times, and take them as the reference image and the target image respectively.
+(4) Text prompt generation.
+In each image pair, the target image is densely captioned with the state-of-the-art large vision language model [1], serving as the text prompt to guide the generation.
+Merits of MC-IDC: Our MC-IDC dataset contains 315k high-quality samples across 10k+ categories and diverse domains, with each sample featuring reference-target image pairs, segmentation masks, and text captions to support ID-preserving, diverse object customization.
+Particularly, it has these merits: (1)Diverse domains and categories: Our MC-IDC dataset contains around 315k samples across 10k+ categories, including human faces, animals, clothes, and tools.
+It features images from diverse domains, such as real-world photos, animations, model-generated content, and movies, enabling general ID understanding.
+Further details are provided in the Supplementary.
+(2)High quality: With an average resolution of 1039$\times$950, our dataset ensures high-quality training and inference for our model.
+(3) Reference-target image pair: Each sample includes a reference image with a segmentation mask, a target image, and a text caption.
+The reference image provides ID information, while the target image and text guide semantic generation.
+The reference and target images depict the same object in different states, sharing the same ID and varying non-ID elements like motion and direction.
+This setup ensures ID preservation while maintaining generation diversity.
 4 Method
-We first outline the overall structure of AnyMaker in Sec.
-3.1.
-Then, we detail the general ID extraction, dual-level ID injection and ID-aware decoupling in Sec.
-3.2, Sec.
-3.3 and Sec.
-3.4 respectively.
-3.1 Overview
-Give a reference image and a text prompt, our goal is to generate images that shares the same ID as the object of interest from the reference image, and modifies non-ID elements such as motions and backgrounds in accordance with the text prompt.
-To this end, we propose AnyMaker, an innovative zero-shot text-to-image customization framework for general objects, as summarized in Fig.
+Overview.
+Given a reference image and text prompt, our goal is to generate an image that retains the identity (ID) of the object in the reference while modifying non-ID elements, like motions and backgrounds, based on the prompt.
+Thus we introduce CustAny, a zero-shot text-to-image customization framework for general objects, as in Fig.
 2.
-Firstly, we utilize the general ID extractor to acquire sufficient ID information from the reference image with a segmentation mask of the interested object.
-Next, we leverage the dual-level ID injection to endow the diffusion model with ID information at the global and local levels respectively, without damaging the text editing ability of the model.
-Additionally, we design a novel ID-aware decoupling module to distinguish the ID information from the coupled non-ID elements in the object representation, for higher ID fidelity and better text editing ability.
-3.2 Genral ID Extraction
-We utilize an ensemble of self-supervised models, consisting of DINOv2 [22] and MAE [23], to extract the representations of general objects in the reference image.
-Previous works [6, 5] targeting at customization tasks in specific domains like human face, usually use a single model to extract object representations.
-However, our experiments demonstrate that one single pre-trained model, such as CLIP [24], DINOv2 [22] or MAE [23], cannot extract sufficient features for high-ID-fidelity generation of general objects, as shown in Fig.
+We first use an ID extractor to capture ID information from the reference image with a segmentation mask.
+Dual-level ID injection is then applied to embed the ID into the diffusion model globally and locally, preserving text-editing capabilities.
+Finally, an ID-aware decoupling module separates ID details from non-ID elements, improving ID fidelity and text-editing accuracy.
+4.1 General ID Extraction
+To extract comprehensive ID information, we use a combination of self-supervised models, DINOv2 [16] and MAE [12], to extract representations for general objects in the reference image.
+Specifically, previous methods, such as [14, 30], which target specific domains like faces, typically use a single model like CLIP [22] to extract object features.
+However, we find that using only one pre-trained model—whether CLIP, DINOv2, or MAE—leaves gaps in capturing features necessary for high-ID-fidelity in general objects (see Fig.
 5 and Tab.
-2.
-Specifically, previous works [4, 5] demonstrate that features extracted by CLIP contain insufficient details, thus being less practical as an ID feature extractor, while DINOv2 is capable of providing desirable details due to its self-supervised contrastive learning strategy.
-However, we find that the features extracted by DINOv2 are color-insensitive because it is trained with augmentations such as ColorJitter.
-Consequently, solely using DINOv2 as an ID extractor would lead to mistakenly generated color.
-MAE, on the other hand, is aware of the color thanks to its reconstruction-related objective.
-Therefore, we leverage DINOv2 and MAE, which compensate for each other, to extract sufficient and comprehensive information to handle the challenging customization task for general objects with various categories and diverse characteristics.
-To align the extracted features and facilitate the subsequent injection, we adopt a two-layer MLP.
-The process of extracting general object representations is as follows,
+2).
+Essentially, while CLIP as in [8, 30] lacks sufficient detail, DINOv2 (despite its strong details due to contrastive learning) is color-insensitive because it is trained with color-augmented techniques like ColorJitter.
+MAE, however, retains color sensitivity due to its reconstruction-based training.
+Thus, combining DINOv2 and MAE allows us to capture both detail and color for more accurate ID extraction.
+We combine DINOv2 and MAE, which complement each other, to capture comprehensive information for the challenging task of customizing diverse general objects.
+To align and prepare these features for injection, we use a two-layer MLP.
+The extraction process for general object representations is as follows,
 $f_{dino}^{C},f_{dino}^{P}=MLP(F_{dino}(I_{ref}\odot M_{ref})),$ (1)
 $f_{mae}^{C},f_{mae}^{P}=MLP(F_{mae}(I_{ref}\odot M_{ref})),$ (2)
-where $\odot$ represents element-wise multiplication, and $f_{dino}^{C},f_{dino}^{P},f_{mae}^{C},f_{mae}^{P}$ represent the class tokens and patch tokens extracted by DINOv2 and MAE respectively.
-$F_{dino}$ and $F_{mae}$ represent the DINOv2 and MAE backbone.
-$I_{ref}$ represents the reference image, and $M_{ref}$ is the segmentation mask of the interested object in the reference image.
+where $\odot$ represents element-wise multiplication; and we have the class tokens and patch tokens $f_{dino}^{C},f_{dino}^{P}$, $f_{mae}^{C},f_{mae}^{P}$ extracted by DINOv2 and MAE respectively.
+$F_{dino}$ and $F_{mae}$ are the DINOv2 and MAE backbone.
+$I_{ref}$ denotes the reference image; and $M_{ref}$ is the segmentation mask of the interested object in the reference image.
 Consequently, tokens derived from our general ID extractor not only offer intricate details from DINOv2, but also have sufficient colour and structural information from MAE.
 4.2 Dual-level ID Injection
-We propose a dual-level ID injection mechanism to inject the global semantic ID information and local fine-grained ID information into the diffusion UNet respectively, ensuring both flexible text editability and high ID fidelity in the customization process for general objects.
+We aim to inject maximum extracted ID information into the diffusion UNet without compromising its text-editing ability.
+Our dual-level ID injection mechanism separately embeds global semantic ID and local fine-grained ID into the UNet, ensuring flexible text edits and high ID fidelity in customizing general objects.
 Global ID injection.
-In text-to-image generation, text prompts are leveraged to guide diffusion models to generate diverse images.
-Nevertheless, diffusion models are incapable of accurately generating images with a specific object ID solely relying on the text description.
-To inject the ID information of the interested objects into the text, while simultaneously maintaining the editing capability of text prompts for other elements such as the background and motions, we design the global ID injection mechanism.
-Specifically, we merge the semantic-level class tokens $(f_{dino}^{C},f_{mae}^{C})$ with the class word in the text embedding, such as "dog" or "cat",
+In text-to-image generation, text prompts guide diffusion models to create diverse images.
+However, diffusion models cannot reliably generate images with a specific object ID based on text alone.
+To embed object ID information into the text while preserving editability for other elements like background and motion, we introduce a global ID injection mechanism.
+Specifically, we merge the semantic-level class tokens $(f_{dino}^{C},f_{mae}^{C})$ with the class word in the text embedding, such as “dog” or “cat”,
 $f_{fuse}^{C}=MLP(Concat(f_{text}^{C},f_{dino}^{C},f_{mae}^{C})),$ (3)
-where $f_{fuse}^{C}$ is the fused class token, and $f_{text}^{C}$ represents the class word embedding of the interested object.
-Then we insert the fused class token $f_{fuse}^{C}$ to the original position of the class word in the text embeddings, obtaining the global-level condition $c_{g}$ with ID-related information of the object of interest, as shown in Fig.
+where $f_{fuse}^{C}$ is fused class token and $f_{text}^{C}$ is class word embedding.
+We insert $f_{fuse}^{C}$ into the class word position in the text embeddings, obtaining the global-level condition $c_{g}$ with ID-related information, as in Fig.
 2.
-Following this, the global-level condition $c_{g}$ engages with the cross-attention module in the UNet, like to standard text-to-image models [10], as follows,
+The global-level condition $c_{g}$ interacts with cross-attention module of UNet, similar to standard text-to-image models [25] as,
 $Z_{g}=Attention(Q_{g},K_{g},V_{g}),$ (4)
@@
 Local ID injection.
-Semantic-level information in the global-level condition $c_{g}$ contains relatively fewer details of the objects, and thus the model cannot handle complex customization tasks for general objects effectively merely relying on global ID injection.
-To supply the model with sufficient ID-related details, we propose the local ID injection mechanism.
-Specifically, we first fuse the extracted patch tokens $f_{dino}^{P}$ and $f_{mae}^{P}$ with a two-layer MLP to densify the contained information to a unified representation $c_{l}$, as follows,
+The global-level condition $c_{g}$ contains limited object details, making it insufficient for complex customization tasks.
+To provide more ID-related information, we propose the local ID injection mechanism, where we fuse the extracted patch tokens $f_{dino}^{P}$ and $f_{mae}^{P}$ using a two-layer MLP to create a unified representation $c_{l}$,
 $c_{l}=MLP(f_{dino}^{P})+MLP(f_{mae}^{P}),$ (5)
@@
 $Z_{l}=Attention(Q_{l},K_{l},V_{l}),$ (6)
-where $Q_{l}=Z_{l}W_{q}$, $K_{l}=c_{l}W_{k}$, $V_{l}=c_{l}W_{v}$ are the query, key, and value of cross-attention module, in which $W_{l}$, $W_{l}$, $W_{l}$ are the weight matrices of the trainable linear projection layers, and $Z_{l}$ is the query feature containing the latent input information.
-Through the dual-level ID injection, we provide the model with sufficient ID-related information, both at the semantic level and the detail level, on the premise of not impairing the text editing ability of the model.
+where $Q_{l}=Z_{l}W_{q}$, $K_{l}=c_{l}W_{k}$, $V_{l}=c_{l}W_{v}$ are the query, key, and value of cross-attention module, in which $W_{l}$, $W_{l}$, $W_{l}$ are the weight matrices of the trainable linear projection layers, and $Z_{l}$ is query feature containing the latent input information.
+By dual-level ID injection, we have the model with sufficient ID-related information at both semantic and detail levels, without impairing its text-editing ability.
 4.3 Training with ID-Aware Decoupling
-The ID-related information injected in the diffusion model is usually entangled with redundant non-ID information such as motion, direction and size.
-Without proper guidance, the model would learn both kinds of information together, which may disrupt the text editing ability of the model and hinder the diversity of generation.
-For instance, in a reference image depicting "a standing person", the posture "standing" may also be extracted and injected into the diffusion model, which could impede a text prompt wishing to generate "a sitting person".
-Consequently, the generated images would closely resemble the provided reference images, thereby compromising the text editing ability and generation diversity.
-To eliminate the negative impact brought by the coupling phenomenon, we take inspiration from previous works [9, 23] to propose an ID-aware decoupling module, which consists of a decoupling branch and a group of losses based on the normal generation branch at the training stage, guiding the model to discern the intrinsic properties of ID information.
+The ID-related information injected into the diffusion model is often mixed with non-ID details like motion, direction, and size.
+Without proper guidance, the model may learn both together, disrupting text editing and limiting generation diversity.
+For example, in a reference image of “a standing person” the posture “standing” could be injected, preventing the generation of “a sitting person” from a text prompt, leading to less diverse outputs.
+To address this, we propose an ID-aware decoupling module, inspired by previous works [5, 12], which includes a decoupling branch and associated losses during training to help the model distinguish ID information from other features.
 Decoupling branch and normal branch.
@@
 Next, we mask out the ID information contained in the image embedding by a trainable feature mask $m_{id}$, and thus the masked image feature $f_{msk}=f_{tar}\odot m_{id}$ contains only non-ID information of the target image, where $\odot$ represents element-wise multiplication.
-Then, we add the masked image feature $f_{msk}$ to the fused class token $f_{fuse}^{C}$ and then feed it to the diffusion UNet for generation following the injection way mentioned in the Sec.
-3.3.
+Then, we add the masked image feature $f_{msk}$ to the fused class token $f_{fuse}^{C}$ and then feed it to the diffusion UNet for generation following the injection way mentioned above.
 Yet in the normal branch, we inject $f_{fuse}^{C}$ to the model without $f_{msk}$.
 The generation process of two branches can be illustrated as follows,
-$f_{msk}\oplus f_{fuse}^{C}\Rightarrow I_{tar},\hskip 14.22636ptf_{fuse}^{C}\Rightarrow I_{tar},\hskip 14.22636ptf_{msk}\perp f_{fuse}^{C},$ (7)
-where $\oplus$ represents element-wise addition and $I_{tar}$ is the target image, and the symbol $\perp$ signifies that the features it connects are independent of each other and have no obvious correlation.
-Through training in the two branches, we push the non-ID information into the masked image feature $f_{msk}$ and only retain the ID-related information in the fused class token $f_{fuse}^{C}$.
+$f_{msk}\oplus f_{fuse}^{C}\Rightarrow I_{tar},\hskip 5.69046ptf_{fuse}^{C}\Rightarrow I_{tar},\hskip 5.69046ptf_{msk}\perp f_{fuse}^{C},$ (7)
+where $\oplus$ denotes element-wise addition, $I_{tar}$ for target image, and $\perp$ indicates that the features on both sides are independent.
+Through training in both branches, we push non-ID information into masked image feature $f_{msk}$ and retain only ID-related information in the fused class token $f_{fuse}^{C}$.
 Training strategy. We use three losses in our training stage to ensure that the ID-aware decoupling module takes effect. Specifically, We utilize the following denoising losses to train each branch,
@@
 $\mathcal{L}_{normal}=\|\epsilon-\epsilon_{\theta}\big{(}x_{t},c_{g},c_{l},t\big{)}\|^{2},$ (9)
-where $t$ is the randomly sampled time step and $x_{t}$ represents the noisy latent of the target image, $\epsilon$ is the ground truth noise and $\epsilon_{\theta}$ is the predicted noise.
-$c_{g}$ and $c_{l}$ represent the global-level and local-level conditions acquired in Sec.
-3.3.
+where $t$ is the randomly sampled time step; and $x_{t}$ represents the noisy latent of the target image; $\epsilon$ is the ground truth noise and $\epsilon_{\theta}$ is the predicted noise.
+$c_{g}$ and $c_{l}$ shall be the global-level and local-level conditions acquired in the dual-level ID injection module.
 Additionally, we design a contrastive loss, ensuring that the masked image feature $f_{msk}$ captures non-ID information and the fused class token $f_{fuse}^{C}$ remains free of it, as below,
@@
 $\mathcal{L}=\alpha_{1}\mathcal{L}_{normal}+\alpha_{2}\mathcal{L}_{decouple}+\alpha_{3}\mathcal{L}_{contrast},$ (11)
-where $\alpha_{1},\alpha_{2},\alpha_{3}$ represent hyperparameters to balance the three losses.
-4 MC-IDC Dataset
-We create the first large-scale general ID dataset, named as Multi-Category ID-Consistent (MC-IDC) dataset, designed for the research of general object customization.
-We introduce the characteristics and construction process of the dataset in Sec.
-4.1 and Sec.
-4.2 respectively.
-We introduce more details of our dataset in the Appendix.
-4.1 Characteristics
-Large-scale dataset.
-Our dataset consists of approximately 315,000 samples in total with more than 10,000 categories, covering various types such as human faces, animals, clothes, human-made tools, etc.
-Adequate training samples and diverse categories endow the model with general ID understanding.
-We introduce the main categories of the dataset and their quantities in the Appendix.
-High quality. The average resolution of the images in our dataset is 1039$\times$950, which is helpful for the high-quality generation of our model in both training and inference stage.
-Reference-target image pair.
-Each sample consists of a reference image, a segmentation mask of the object of interest in the reference image, a target image, and a text caption of the target image.
-During training, the reference image with its segmentation mask provides the model with ID information, the text caption of the target image offers semantic-level guidance for generation, and the target image serves as the ground truth.
-In each sample, the reference image and the target image depict the same object in different states, which means that each reference-target image pair shares the same ID of the object of interest and has different non-ID elements such as motions and directions.
-Training with reference-target image pairs endows the model with the ID processing ability while not damaging the diversity of generation.
-4.2 Construction Pipeline
-The construction pipeline of MC-IDC consists of four steps: data collection, instance detection and segmentation, image pair generation, and text prompt generation.
-Data collection.
-MC-IDC dataset contains diverse sources including web-crawled images, movies and publicly available datasets.
-To facilitate the subsequent generation of image pairs, most of the data we collect are video datas or multi-view datas.
-To ensure the high quality of our dataset, we delete the images with a resolution less than 300$\times$300.
-Instance detection and segmentation.
-For some publicly available datasets or web-crawled images which do not provide segmentation annotations, we use the advanced instance segmentation model [25] for instance detection and segmentation.
-For movie dataset, we extract the frames with 10fps and do object tracking with  [26], aiming to establish the ID connection across frames.
-Image pair generation.
-For most data samples which are from a clip of video or a group of multi-view data, we randomly select two frames containing the same object and perform a random crop around the object to form the reference-target image pair.
-For the other data samples from the single image dataset, we perform random data augmentations on the single image two times, and take them as the reference image and the target image respectively.
-Text prompt generation. In each image pair, the target image is densely captioned with the state-of-the-art large vision language model [27], serving as the text prompt to guide the generation.
-For more details about the MC-IDC, please refer to the Appendix.
+where hyperparameters $\alpha_{1},\alpha_{2},\alpha_{3}$ are coefficients which are set to 2.0, 1.0, 0.5 respectively.
 Table 1: Quantitative comparison among zero-shot object customization methods in general domain and two specific popular domains namely human customization and virtual try-on.
@@
 5 Experiment
-5.1 Setup
 Implementation details.
-To be compatible with open-community as Civitai.com, we employ SD1.5 [28], the most used foundational model in the public, as the backbone model in our experiments.
-We use the multi-scale training mode proposed in [1] considering the resolutions of images from different sources are diverse.
-We employ the same CLIP image encoder as in [28] in the ID-aware decoupling module.
-We use MAE [23] and DINOv2 [22] for extracting ID features.
-During training, we set the learning rate as 1e-5 with a batch size of 32.
-We conduct the training process about 6 epochs with 32 V100 GPUs, consuming about 30 hours.
-We adopt dropout with a ratio of 0.1 for the global-level condition $c_{g}$ and the local-level condition $c_{l}$ respectively.
-During inference, we utilize 50 steps denoising process for generation and set the classifier-free guidance scale as 7.
+We use SD1.5 as the backbone model for compatibility with open communities like Civitai.com, and apply the multi-scale training mode from [21] to handle varied image resolutions.
+The ID-aware decoupling module uses the CLIP encoder from [26], while MAE-ViT-h/14 and DINOv2-ViT-g/14-reg4 are used to extract ID features.
+Training involves a 1e-5 learning rate, batch size of 32, and 6 epochs on 32 V100 GPUs, taking about 30 hours.
+During inference, we perform 50 denoising steps and set the classifier-free guidance scale to 7.
 Competitors.
-To demonstrate the generality of our AnyMaker, we not only make a comparison with previous works on general object customization, but also compare it with task-specialized methods in two popular specific customization domains, that is, human customization and virtual try-on.
-Specifically, we compare AnyMaker with IP-Adapter [7] on zero-shot general object customization.
-On human customization, we further choose human specialized method PhotoMaker [6] and InstantID [5] for comparison.
-On virtual try-on, we conduct a comparison with the recent state-of-the-art MagicClothing  [21].
-It is worth noting that all the methods mentioned above are zero-shot methods and are capable of customizing images based solely on one single reference image.
-Therefore, in this paper, we do not compare object-specific methods such as DreamBooth [8] and DisenBooth [9] which typically require 3 to 5 reference images and additional fine-tuning.
+To demonstrate CustAny’s generality, we compare it with both general object customization methods and task-specific methods in human customization and virtual try-on.
+We compare CustAny with IP-Adapter [33] for general object customization, PhotoMaker [14] and InstantID [30] for human customization, and MagicClothing [7] for virtual try-on.
+All methods are zero-shot and use a single reference image, so we exclude object-specific methods like DreamBooth [27] and DisenBooth [5] that require multiple reference images and fine-tuning.
 Evaluation dataset.
-Our evaluation dataset contains 1000 text-image samples, including general objects, human data, and virtual try-on data with a ratio of 4:3:3.
-All samples in the evaluation dataset are randomly selected from the MC-IDC and do not appear during the training stage.
+Our evaluation dataset consists of 1,000 text-image samples, covering general objects, human data, and virtual try-on data in a 4:3:3 ratio.
+None of these samples are included in the training set.
 Evaluation metrics.
-Following PhotoMaker [6], we utilize DINO-i [22] and CLIP-i [29] to measure the ID fidelity and use CLIP-t to measure the prompt fidelity.
+As PhotoMaker [14], we utilize DINO-i [16] and CLIP-i [23] to measure the ID fidelity and use CLIP-t to measure the prompt fidelity.
 We leverage FID [13] to assess the generation quality.
 For human customization, we additionally calculate the face similarity (FaceSim) with FaceNet [28] as commonly done in [14, 30].
-Furthermore, we design a novel metric DiverSim-i, which calculates the average DINO similarity among the images generated according to diverse text prompts containing different scenarios.
-The lower the value of DiverSim-i, the stronger the ability of the model to generate diverse images in line with text prompts.
-We introduce more details about the experimental setup in the Appendix.
-Figure 3: Qualitative comparison on zero-shot general object customization and two specific popular tasks namely human customization and virtual try-on.
-The AnyMaker exhibits great ID-preserving ability with better text controls and more diverse generations on both general objects and specialized domains.
-5.2 Quantative Analysis
-Our AnyMaker outperforms previous works on general object customization for all metrics, and achieve comparable or better performance in contrast with specific methods tailored for specialized tasks such as human customization and virtual try-on, as shown in Tab.
+Further, we introduce DiverSim-i, a novel metric that measures the average DINO similarity across images generated from diverse text prompts.
+A lower DiverSim-i value indicates a stronger model ability to generate diverse images matching the prompts.
+Details are in Supplementary.
+Figure 3: Qualitative results on general domains and two specific specific domains: human customization and virtual try-on.
+CustAny exhibits great ID-preserving ability with better text controls and more diverse generations on both general objects and specialized domains.
+Figure 4: Results: the same reference in different text prompts.
+5.1 Quantitative and Qualitative Analysis
+Quantitative results.
+Our CustAny outperforms previous works on general object customization for all metrics, and achieve comparable or better performance in contrast with specific methods tailored for specialized tasks such as human customization and virtual try-on, as shown in Tab.
 1.
@@
 with decoupling 47.09 82.16 65.13
-5.3 Qualitative Analysis
-The AnyMaker exhibits outstanding capabilities of high-quality customization for general objects, and even beat task-specialized methods in the specific domains, such as human customization and virtual try-on, as shown in Fig.
+Qualitative analysis.
+The CustAny exhibits outstanding capabilities of high-quality customization for general objects, and even beat task-specialized methods in the specific domains, such as human customization and virtual try-on, as shown in Fig.
 3.
+We also show the visual results of the same reference picture under different text prompts in Fig.
+4.
 Our method is capable of generating high-ID-fidelity images given one single reference image, and simultaneously editing non-ID elements like posture or background in accordance with text prompts, which endows the model with the ability to generate diverse images, such as various scenarios and different postures.
 Furthermore, benefiting from the ID-aware decoupling module, our model is capable of maintaining the ID fidelity of objects while diversifying the non-ID elements such as motions and directions compared to the reference images, even without corresponding guidance of text prompts.
-Figure 4: Ablation study on different ID extraction methods in the quantitative perspective. Our ID extractor has both the rich details from DINOv2 and the color information from MAE.
-Figure 5: Ablation study on whether to add the ID-aware decoupling module in the quantitative perspective.
+We show more visual results in Supplementary.
+Figure 5: Ablation on different ID extraction methods in the quantitative perspective. Our ID extractor has both the rich details from DINOv2 and the color information from MAE.
+Figure 6: Ablation on whether to add the ID-aware decoupling module in the quantitative perspective.
 The model with the ID-aware decoupling module demonstrates stronger text control ability, as shown by the blue text.
-5.4 Ablation Study
 Effectiveness of general ID extraction.
@@
 As in Fig.
-4, merely using DINOv2 as the extractor fails to maintain color consistency, while solely using MAE as the extractor lacks sufficient details.
+5, merely using DINOv2 as the extractor fails to maintain color consistency, while solely using MAE as the extractor lacks sufficient details.
 In contrast, our method combines the advantages of both DINOv2 and MAE, which enables our model to customize images with rich details and the consistent color of the object in reference images.
@@
 As shown in Tab.
-3, by combining global and local injection, AnyMaker achieves the best generation results both in quality as measured by FID and ID-fidelity as measured by CLIP-i and DINO-i.
+3, by combining global and local injection, CustAny achieves the best generation results both in quality as measured by FID and ID-fidelity as measured by CLIP-i and DINO-i.
 Benefits of ID-aware decoupling.
 We conduct a comparative experiment on whether to add the ID decoupling module to the model during training, aiming to verify its effectiveness.
-As in Tab.4, the model trained with the ID-aware decoupling module achieves higher ID-fidelity scores in terms of CLIP-i and DINO-i, which means that the decoupling module can help the model to discern ID information embedded in object representations, thereby generating results with better ID consistency.
-Further, we visualize the results by models trained with and without decoupling in Fig.5.
+As in Tab.
+4, the model trained with the ID-aware decoupling module achieves higher ID-fidelity scores in terms of CLIP-i and DINO-i, which means that the decoupling module can help the model to discern ID information embedded in object representations, thereby generating results with better ID consistency.
+Further, we visualize the results by models trained with and without decoupling in Fig.
+6.
 The model trained with decoupling exhibit enhanced capabilities in distinguishing ID information from non-ID elements such as motions (open or shut the mouth) and directions (front-facing or side-facing) of objects of interest.
 The enhanced discrimination ability allows the model to mitigate the influence of non-ID information during generation, thereby better preserving the text editing capabilities.
+5.2 Additional Qualitative Applications
+Text-image ID mixing. If the category of the interested object in the text prompt and that in the reference image is not the same, our CustAny can merge the two and form a new ID, as shown in Fig. 7
+Story generation.
+Our CustAny can generate diverse images under the guidance of text prompts, while maintaining the same identity as the object of interest in the reference image, thereby enabling the creation of a cohesive narrative, as shown in Fig.
+8.
+Figure 7: Applications of CustAny on text-image ID mixing, such as merging the text ”dog” with the image ”tiger”.
+Figure 8: Applications of CustAny on story generation. Given continuous text prompts and one image example, CustAny can generate corresponding stories.
 6 Conclusion
-We introduce AnyMaker, a zero-shot text-to-image framework for general object customization, which has seldom been studied in previous works.
-Our approach integrates novel general ID extraction, dual-level ID injection, and ID-aware decoupling module, endowing the model with high ID fidelity without damaging the text editing ability.
-Experiments demonstrate that our method excels in general object customization solely with one single reference image, and even outperforms task-specialized methods in some popular specific domains.
-Furthermore, we construct the first large-scale general ID dataset, MC-IDC, to set a new benchmark and promote the future research for object customization.
-Appendix A Appendix
-A.1 Additional Details about MC-IDC
+We introduce CustAny, a zero-shot text-to-image framework for general object customization, integrating general ID extraction, dual-level ID injection, and an ID-aware decoupling module.
+CustAny achieves high ID fidelity while preserving text editing abilities and outperforms specialized methods in certain domains.
+Additionally, we create the MC-IDC dataset, the first large-scale general ID dataset, promote the research for object customization.
+[…]
+Supplementary Material
+7 Additional Details about MC-IDC
+Main categories. We record several main categories that appear most frequently in MC-IDC, as shown in Tab. 5.
 Data sources.
@@
 We detail the statistical information about various data sources in Tab.
-A1.
-Main categories. We record several main categories that appear most frequently in MC-IDC, as shown in Tab. A2.
-Figure A1: Illustration of our general ID dataset, MC-IDC.
-In each sample, the reference image with the object mask provides ID information, the text prompt offers semantic-level guidance for generation, and the target image serves as the ground truth.
-Dataset illustration. We present the overall illustration of the MC-IDC as shown in Fig. A1.
-Table A1: Details about data sources of MC-IDC.
-Source Dataset Type Image pair numbers
-Public datasets HumanFace [32] Video 55,830
-VOS [33] Video 55,823
-VIPSEG [34] Video 27,983
-MVImgNet [35] Multi-view image 53,909
-VITON [36] Multi-view image 20,000
-LVIS [37] Single image 8,003
-Web-crawled images - Single image 55,829
-Movies - Video 38,405
-Table A2: Sample numbers of main categories in MC-IDC.
+6.
+Table 5: Sample numbers of main categories in MC-IDC.
 Categories Images
@@
 dog 1,820
-A.2 Additional Details about Experiment Setup
+Table 6: Details about data sources of MC-IDC.
+Source Dataset Type Image pair numbers
+Public datasets HumanFace [2] Video 55,830
+VOS [32] Video 55,823
+VIPSEG [15] Video 27,983
+MVImgNet [34] Multi-view image 53,909
+VITON [11] Multi-view image 20,000
+LVIS [10] Single image 8,003
+Web-crawled images - Single image 55,829
+Movies - Video 38,405
+8 Additional Details about Experiment Setup
 Categories in evaluation dataset.
@@
 We summarize the 50 categories in Tab.
-A3.
-Table A3: Categories of general objects in the evaluation dataset.
+7.
+Table 7: Categories of general objects in the evaluation dataset.
 Winter melon Cabbage Vessel Pillow Screw driver
@@
 We record the text prompts in Tab.
-A4.
-Table A4: Text prompts for calculating DiverSim-i.
+8.
+Table 8: Text prompts for calculating DiverSim-i.
 Scenarios Text prompts
-Snow Original text prompt + "The scene of the picture is in the snow."
-Original text prompt + "The background of the picture is in the snow."
-Grass Original text prompt + "The scene of the picture is on the grass."
-Original text prompt + "The background of the picture is on the grass."
-Beach Original text prompt + "The scene of the picture is on the beach."
-Original text prompt + "The background of the picture is on the beach."
-Jungle Original text prompt + "The scene of the picture is in the jungle."
-Original text prompt + "The background of the picture is in the jungle."
-Eiffel Tower Original text prompt + "The scene of the picture is beside the Eiffel Tower."
-Original text prompt + "The background of the picture is beside the Eiffel Tower."
-A.3 More Visual Results
-Our AnyMaker exhibits outstanding performance on various applications, such as general object customization in Fig.
-A2, virtual try-on in Fig.
-A3, Text-Image ID Mixing in Fig.
-A4, ID-consistent generation in Fig.
-A5, and story generation in Fig.
-A6
-Figure A2: Additional visual results of general object customization.
-Our AnyMaker is capable of generating high-ID-fidelity images given one single reference image, and simultaneously editing non-ID elements like posture or background in accordance with text prompts.
-Figure A3: Additional visual results of virtual try-on. Given a piece of clothing, the AnyMaker can generate images of the clothing worn on a person.
-Figure A4: Additional visual results of text-image ID mixing.
-If the category of the interested object in the text prompt and that in the reference image is not the same, our AnyMaker can merge the two and form a new ID.
-Figure A5: Additional visual results of ID-consistent generation. The AnyMaker can generate multiple ID-consistent images with diverse non-ID elements such as motions and orientations.
-Figure A6: Additional visual results of story generation.
-Our AnyMaker can generate diverse images under the guidance of text prompts, while maintaining the same identity as the object of interest in the reference image, thereby enabling the creation of a cohesive narrative.
+Snow Original text prompt + ”The scene of the picture is in the snow.”
+Original text prompt + ”The background of the picture is in the snow.”
+Grass Original text prompt + ”The scene of the picture is on the grass.”
+Original text prompt + ”The background of the picture is on the grass.”
+Beach Original text prompt + ”The scene of the picture is on the beach.”
+Original text prompt + ”The background of the picture is on the beach.”
+Jungle Original text prompt + ”The scene of the picture is in the jungle.”
+Original text prompt + ”The background of the picture is in the jungle.”
+Eiffel Tower Original text prompt + ”The scene of the picture is beside the Eiffel Tower.”
+Original text prompt + ”The background of the picture is beside the Eiffel Tower.”
+9 More Visual Results
+Our CustAny exhibits outstanding performance on various applications, such as story generation in Fig.
+9, virtual try-on in Fig.
+10, and ID-consistent generation in Fig.
+11.
+Our CustAny can ensure both the ID fidelity and the generating diversity simultaneously.
+Figure 9: Additional visual results of story generation.
+Our CustAny can generate diverse images under the guidance of text prompts, while maintaining the same identity as the object of interest in the reference image, thereby enabling the creation of a cohesive narrative.
+Figure 10: Additional visual results of virtual try-on. Given a piece of clothing, the CustAny can generate images of the clothing worn on a person.
+Figure 11: Additional visual results of ID-consistent generation. The CustAny can generate multiple ID-consistent images with diverse non-ID elements such as motions and orientations.